Exercise 22-15 The first audit of the books of Crane Company was made for the year ended December 31, 2018. In examining the books, the auditor found that certain items had been overlooked or incorrectly handled in the last 3 years. These items are: 

  1. At the beginning of 2016, the company purchased a machine for $528,000 (salvage value of $52,800) that had a useful life of 6 years. The bookkeeper used straight-line depreciation but failed to deduct the salvage value in computing the depreciation base for the 3 years.

  2. 2. At the end of 2017, the company failed to accrue sales salaries of $49,000 which was paid in 2018 and was debited to Salaries and Wages Expense.

  3. 3. A tax lawsuit that involved the year 2016 was settled late in 2018. It was determined that the company owed an additional $81,000 in taxes related to 2016. The company did not record a liability in 2016 or 2017 because the possibility of loss was considered remote, and charged the $81,000 to a loss account in 2018.

  4. 4. Crane Company purchased a copyright from another company early in 2016 for $52,000. Crane had not amortized the copyright because its value had not diminished. The copyright has a useful life at purchase of 20 years.

  5. 5. In 2018, the company wrote off $94,000 of inventory considered to be obsolete; this loss was charged directly to Retained Earnings. Prepare the journal entries necessary in 2018 to correct the books, assuming that the books have not been closed. Disregard effects of corrections on income tax. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter o for the amounts. Round answers to O decimal places, e.g, 1250.) No. Account Titles and Explanation Debit Credit 1. 2.
Answer #1

Solution:

Particulars

Debit

Credit

Accumulated Depreciation—Machinery

44,000

Depreciation Expense

17,600

Depreciation Expense

26,400

Retained Earnings

49,000

Sales Salaries Expense

49,000

No entry necessary

Amortization Expense—Copyright (52,000/20years)

2,600

Retained Earnings ($2600 * 2 months)

5,200

Copyright

7,800

Loss on Write-down of Inventories

94,000

(or Cost of Goods Sold)

94,000

Retained Earnings

Working:

2016-17

2018

528000*1/6*2 = 176,000

52800*2 = 105600

79200 * 2 years = 158400

(528000 - 52800) / 6 years = 79,200

17600

26400